OpenStudy (anonymous):

What is a polynomial function in standard form with zeroes 1, 2, –3, and –3?

OpenStudy (anonymous):

ou can't have a quadratic function with four zeros. This will be a quartic function (4th degree polynomial function) zeros 3, 1, 2, −3 -----> polynomial has roots (x−3), (x−1), (x−2), (x+3) f(x) = a (x−3) (x−1) (x−2) (x+3) f(x) = a (x−3)(x+3) * (x−1)(x−2) f(x) = a (x² − 9) (x² − 3x + 2) f(x) = a (x⁴ − 3x³ + 2x² − 9x² + 27x − 18) f(x) = a (x⁴ − 3x³ − 7x² + 27x − 18)

OpenStudy (anonymous):

you*

OpenStudy (anonymous):

if you're flipping the signs wouldn't it then be (x+3) (x−1) (x−2) (x+3)?

OpenStudy (anonymous):

would it?

OpenStudy (anonymous):

What have you learned?

OpenStudy (anonymous):

Not a whole lot. really confused

OpenStudy (anonymous):

Are you homeschooled?

hartnn (hartnn):

oh, it would be (x+3) (x−1) (x−2) (x+3)

OpenStudy (anonymous):

hartnn is good at it

hartnn (hartnn):

with your comment, even you seem to be good..

OpenStudy (anonymous):

pshhhhhhh, www.google.com

OpenStudy (anonymous):

works quite well

OpenStudy (anonymous):

google doesn't work as well as you think.

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/complex/Complex_Numbers.faq.question.117426.html this is kind of like a mini teacher

OpenStudy (anonymous):

not really helping me, sorry

hartnn (hartnn):

f(x) = (x−3) (x−1) (x−2) (x-3) now its just algebra...

OpenStudy (anonymous):

its too much like geometry. I just don't grasps the concepts very well

OpenStudy (anonymous):

still need some help. not askin for and answer just a nudge in the right direction

hartnn (hartnn):

f(x) = (x−3) (x−1) (x−2) (x-3) can u foil ? f(x) = (x^2-6x+9) (x−1) (x−2)

OpenStudy (anonymous):

yes I can foil. though should it be x+3 since they're negative to begin with. that be what is confusing me. and not sure how to foil with three numbers

hartnn (hartnn):

ok, if a is a zero, then x-a is the factor if -3 is a zero, then x+3 is the factor

hartnn (hartnn):

it should be then, f(x) = (x+3) (x−1) (x−2) (x+3) f(x) = (x^2+6x+9) (x−1) (x−2)

hartnn (hartnn):

(x−1) (x−2)= ... ?

OpenStudy (anonymous):

x^2 - 3x + 2

hartnn (hartnn):

right, so u have f(x) = (x^2+6x+9) (x^2-3x+2) so u have difficulty foiling this ?

OpenStudy (anonymous):

its this step I'm having trouble with. because I ended up with x^4 + 7x^2 + 108

hartnn (hartnn):

ok, first multiply x^2 (of (x^2+6x+9) ) with (x^2-3x+2)

hartnn (hartnn):

what u get ?

OpenStudy (anonymous):

x^4

hartnn (hartnn):

just x^4 ? there will be 3 terms , right ?

OpenStudy (anonymous):

x^4 - 3x^3 + 2x^2?

hartnn (hartnn):

yes, thats correct now multiply 6x (of (x^2+6x+9) ) with (x^2-3x+2)

OpenStudy (anonymous):

6x^3 - 18x^2 + 12x

hartnn (hartnn):

yes, thats correct now multiply 9 (of (x^2+6x+9) ) with (x^2-3x+2)

OpenStudy (anonymous):

9x^2 - 27x + 18

hartnn (hartnn):

now add all the 3 results, and that will be your answer.

OpenStudy (anonymous):

oh okay. Thank you =^..^=

hartnn (hartnn):

if u want me to verify, then write what final answer u got...

OpenStudy (anonymous):

x^4 + 3x^3 - 7x^2 - 15x + 8

hartnn (hartnn):

i get 18 in the end

hartnn (hartnn):

+18, other terms are correct

OpenStudy (anonymous):

my bad its a typo. I meant to put 18

hartnn (hartnn):

then its correct